The doctor orders Dopamine 400 mg in 250mL of D3W to be infused at 2mcg/kg/min. The patient weighs 150 lbs. The drop factor is 15 gtts/mL. Calculate the flow rate in mL/hr.

a. 5.1 mL/hr

b. 5.11 mL/hr

c. 5 mL/hr

d. 5.0 mL/hr

e. 5 L/hr
Answer #1

Patient weight - 150 lbs = 68 kgs

Dopamine 2 mcg/ kg/ min = 68 \times 2 = 136 mcg/ min

For 1 hour 136 \times 60 = 8160 mcg

8160 mcg = 8.16 mg

Required medication = required dose \times available volume \div available dose

                                  = 8.16 \times 250 \div 400

                                  = 5.1 ml

Flow rate = total volume \div infusion time (hr)

               = 5.1 \div 1

              = 5.1 ml per hour

Option a is correct answer.